A principal of $4300 is invested at 9% interest, compounded annually. How much will the investment be worth after 7 years?
Answer by mananth(16946)

Find Amount
Principal P = 4300
Amount= A
years=t 7.00
compounded 1 times a year n
Rate = 9.00% 0.09
Amount = P*((n+r)/n)^n*t

Amount = = 4300 *( 1 + 0.09 )^ 1 * 7.00
Amount = 4300 *( 1 + 0.09 )^ 7
4300 *( 1.09 )^ 7.00
Amount = $7860.57